The large hand of a clock is 42cm long.How many centimetres does it extremity move in 20 minutes?

Grade:12

2 Answers

sushant singh
12 years ago

hi gurwinder

20 minutes means 4 x 30o = 120o

so if you calculate the sector circumference , you will get the answer as

2 x Pi x R (120/360) = 2 x 22 /7  x 42 x 1/3  = 88 cm Ans.

n o
12 years ago

now radius(R)= 42cm

cicum(C)=2pi*42=264cm

in 1 min...a distance of '1C' or 264 cm is covered

so in 20mins.... 20C=20*264=5280cm is covered..

plz approve by clicking 'yes'